Tuesday, 25 August

Other Events

\underline{14:00 – 15:45 MA 004}

Mathematics in the Third Reich and Racial and Political Persecution A forum with lectures by Joel Lebowitz (Rutgers University, USA): “Victims, Oppressors, Activists, and Bystanders: Scientists’ Response to Racial and Political Persecution” and by Herbert Mehrtens (Universität Braunschweig, Germany): “Mathematics and Mathematicians in Nazi Germany. History and Memory” and a discussion chaired by Jochen Brüning (HU Berlin, Germany).

\underline{16:00 Urania}

Mathematics and Art The best entries by high-school students will be honored, and some math videos will be shown.

\underline{17:30 Urania}

Gero von Randow, Hamburg, Germany Geisterfahrer des Journalismus — Wie kommt die Mathematik in die Zeitung? (in German)

\underline{18:00 Urania}

Movie Seeing is Believing / Hotel Hilbert

\underline{19:30 Urania}

Walter Schachermayer, Wien, Austria The Role of Mathematics in Financial Markets. The lecture is followed by a round-table discussion involving R. Dantschke, H. Föllmer (chair), A. Gruber, G. Stahl.

\underline{20:00 Urania}

VideoMath Festival The “VideoMath Reel”, the 24 best contributions to the VideoMath competition, and the Award Ceremony

ICM’98 – Program  1